OpenStudy (anonymous):

a solution of Fe(ii) chloride is prepared by dissolving 0.1g og FeCl2(s) in 1kg of water.what are the concentrations in molal of Fe 2+(aq) and Cl- (aq) in the resulting solution?

OpenStudy (aaronq):

convert the mass given to moles, write a reaction. Use the stoichiometric coefficients to find how much (in moles) of each product is produced. e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. Convert to molality: \(m=\dfrac{n_{solute}}{kg_{solvent}}\)

OpenStudy (anonymous):

are you there?

OpenStudy (aaronq):

?